An-8 sided die is rolled and a coin is flipped. Christine gets to pick the movie that her and her boyfriend go to if the die is rolled as less than a 4 and the coin is flipped tail. if this does not occur, her boyfriend picks the movie. Who likely came up with this game? Why?
Answer #1

P(Christine wins) = P(die is rolled as less than a 4 and the coin is flipped tail)

= P(die rolling 1, 2 or 3) x P(tail on coin toss)

= 3/8 x 1/2

= 0.1875

P(Christine losing) = 1 - 0.1875

= 0.8125

The probability of Christine getting to pick the movie is low according to this game. So, Christine's boyfriend must have come up with this game